A father is 20yrs older than his son. If the sum of their ages is 40yrs, how old is the son now?
Answer by ReadingBoosters(3246)

f = s + 20
f + s = 40
...
Substitute the first equation into the second
s+20 + s = 40
Combine like terms
2s + 20 = 40
2s = 40 - 20
2s = 20
s = 10
f = s + 20 = 10 + 20 = 30
...
Father is 30, Son is 10
……………….
Have more questions? Get them all answered at once.
HomeworkHelpers@readingboosters.com
Understand your homework, prepare for upcoming tests, improve your scores.
www.MyHomeworkAnswers.com